OpenStudy (amistre64):

What is the indefinite integration of:\[\int x4^x(3+ln(x^6)+3xln(4)ln(x))dx\]

OpenStudy (anonymous):

i think this proplem is based on integration by parts

OpenStudy (amistre64):

i spose it could be based on that. But this is one of my own creations :)

OpenStudy (amistre64):

it even broke the wolf

OpenStudy (amistre64):

@FoolForMath hows this for a problem of the day?

OpenStudy (anonymous):

soft

OpenStudy (anonymous):

Looks scary.

OpenStudy (amistre64):

that it does, latex makes everything looks scary tho

OpenStudy (anonymous):

I am going to scan my solution this is fun lol

OpenStudy (anonymous):

how do you integrate x*lnx*4^x lol @amistre64

OpenStudy (amistre64):

:) hint: expand the integrand out and simplify it all before trying to integrate

OpenStudy (anonymous):

yeah i did that i got 3 parts i am doing the second part never encountered that integral before lol

OpenStudy (anonymous):

i think i got it nvm lol

OpenStudy (amistre64):

the 2nd one becomes x 4^x ln(x^6) 6x 4^x ln(x)

OpenStudy (anonymous):

yupp i got that just integrating my life away now lol

OpenStudy (amistre64):

dont integrate them separately; look at it as a whole and try to think of a simpler idea that unites them

OpenStudy (amistre64):

the ln(4) should be a dead give away to my evilness lol

OpenStudy (anonymous):

yeah i saw that lol :P

OpenStudy (anonymous):

almost finished do i get a hint in how to integrate lnx*4^x

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

i take it that is from the last term?

OpenStudy (anonymous):

yeahh

OpenStudy (amistre64):

\[x4^x3xln(4)ln(x)\] is rather messy; but lets organize it better \[3x^2\ ln(x)\ 4^xln(4)\] \[[3x^2]\ [ln(x)]\ [4^xln(4)]\]

OpenStudy (anonymous):

yeah so i did this by integration by parts and made the first 2 terms u and the last term dv and then my du=2xlnx +x and dv=4^x

OpenStudy (amistre64):

if we look at the exapnsion; notice the each term has similar parts\[3x4^x+6x4^xln(x)+3x^2ln(x)4^xln(4)\] and recall the product rule:\[\frac{d}{dx}fg=f'g+fg'\] this expands to more than 2 factors right? im not sure what the by parts method would bring us, but its messy

OpenStudy (anonymous):

everything is done i just neeed to know how to integrate lnx*4^x

OpenStudy (anonymous):

do you know how to integrate lnx*4^x because i cant think of any way to integrate this in my past knowloedge lol :P if not i am just going to post my solution

OpenStudy (amistre64):

i cant think of a way to integrate ln(x) 4^x unless we try parts ln(x) 4^x ln(4) - {S} 4^x ln(4)/x dx that doent seem to play nice does it

OpenStudy (amistre64):

i did that backewards; i tend to derive the integrating part :/

OpenStudy (anonymous):

nope then i get (4^x)/x how do we integrate that ?

OpenStudy (amistre64):

by parts isnt a gaurentee that you get a nicer or easier function to play with is it?

OpenStudy (amistre64):

int by parts is a technique, not a rule

OpenStudy (anonymous):

yeah its a technique

OpenStudy (amistre64):

as such, we mmay end up with a harder integration to play with and have to back track or go another route

OpenStudy (amistre64):

the other route to me is to see the problem in a different manner

OpenStudy (amistre64):

assume that the results follow from something that is easier known 3 terms with like parts COULD be a result of:\[\frac{d}{dx}fgh=f'gh+fg'h+fgh'\] and then look for a way that this can match

OpenStudy (amistre64):

we can even format that to a similar by parts formula \[fgh=\int f'gh+\int fg'h+\int fgh'\] \[\int fgh'=fgh-\int f'gh-\int fg'h\]

OpenStudy (anonymous):

alright time to post it lol only one integral couldnt be solved lol other than that its okay @amistre64

OpenStudy (amistre64):

that was a good effort at least :) i devised this from the concept that: \[\frac{d}{dx}3x^2\ 4^x\ ln(x)=3x'^2\ 4^x\ ln(x)+3x^2\ 4'^x\ ln(x)+3x^2\ 4^x\ ln'(x)\] \[\frac{d}{dx}3x^2\ 4^x\ ln(x)=6x\ 4^x\ ln(x)+3x^2\ 4^xln(4)\ ln(x)+3x\ 4^x\] \[\frac{d}{dx}3x^2\ 4^x\ ln(x)=x4^x(3+ln(x^6)+3x\ ln(4)\ ln(x))\]

OpenStudy (anonymous):

@Aka_966